OpenStudy (anonymous):

can someone plz check my answers What are the zeros of the polynomial function: f(x) = (x + 3)(x – 1)(x – 5)? –3, 1, 5 –3, –1, –5 3, 1, –5 3, –1, 5--->MY ANSWER What is the remainder when (x3 – 59) ÷ (x – 4) ? 5 9 --->MY ANSWER What are the zeros of f(x) = x3 – 4x2 – 7x + 10 ? –2, 1, 5 2, –1, –5 –2, 1, –5 --->MY ANSWER –2, –1, –5 1 6

OpenStudy (anonymous):

Not quite ....There are only 3 zeros for the first function. You listed too many.

OpenStudy (anonymous):

THe remainder is not 59

OpenStudy (anonymous):

so the first one i need to resolve

OpenStudy (anonymous):

Again in the last function you have too many zeros.

OpenStudy (anonymous):

ok thanks fot letting me know

OpenStudy (anonymous):

The first one you got mixed up, set the factors (x + 3)(x – 1)(x – 5) all equal to 0 respectively and you get x+3=0 x= (-3) x-1=0 x= (1) x+5=0 x= (-5) Basically a factor is just : (x-a) where a is postitive and (x+a) is just equal to (x-(-a))

OpenStudy (anonymous):

Yes the zeros of a function are only the values of x that will make the function equal zero.

OpenStudy (anonymous):

So Since f(x) = (x+3) (x-1) (x-5) you are looking for what x s make (x+3) (x-1) (x-5) =0

OpenStudy (anonymous):

of course that would be either when x +3 = 0 or when x-1 =0 or when x-5 =0.

OpenStudy (anonymous):

oh ok i see now i didnt set them equal to zero

OpenStudy (anonymous):

So what are your three zeros @countrycuties16

OpenStudy (anonymous):

-3,1,5

OpenStudy (anonymous):

Bingo

OpenStudy (anonymous):

For the second question ... do you know how to do synthetic division?

OpenStudy (anonymous):

i am not good at it i really didnt understand it my teacher refuses to explain it to me

OpenStudy (anonymous):

ok so for synthetic division you list each of the coefficients of each term (do you know what a coefficient is?)

OpenStudy (anonymous):

thee coefficent is x^3?

OpenStudy (anonymous):

no the coefficient is the number you multiply the variables by ... in general it is the number in front of the variable so for example... 6x^2-2t +y = 16 6 is the coefficient of x^2, -2 is the coefficient of t and 1 is the coefficient of y.

OpenStudy (anonymous):

Notice that I included a negative if it was a substraction like -2 for the t term

OpenStudy (anonymous):

oh ok i see so it will be -4

OpenStudy (anonymous):

well yes that is one but hold on.

OpenStudy (anonymous):

so you take the polynomial that you are dividing (not the one that you are dividing by)...so in this case \[(x^3-59) \div (x-4)\] Which one are you dividing?

OpenStudy (anonymous):

(x^3-59)

OpenStudy (anonymous):

right! Now...list all of the coefficients of each term of that polynomial but keep in mind if a term is missing that the coefficient is then 0... so go ahead and list them

OpenStudy (anonymous):

so i did that and i got a remaider of 5

OpenStudy (anonymous):

There you go! Now you have it.

OpenStudy (anonymous):

thanks so much is my last on ok???

OpenStudy (anonymous):

one*

OpenStudy (anonymous):

Now for the last part you have to do a little bit of everything. First list all possible zeros which you do by taking the constant term and listing all of its factors (both positive and negative and call that set P.

OpenStudy (anonymous):

So for example if the function is f(x) = 5x^3 - 4x^2 +5x -8. Then your set P = {1,2,4,8,-1,-2,-4, -8}

OpenStudy (anonymous):

Then find all the factors of the coefficient of the highest degreed term and call those set Q. In my example that would be Q= {1,5,-1,-5}

OpenStudy (anonymous):

Are you following me?

OpenStudy (anonymous):

yes i am i am starting to understand

OpenStudy (anonymous):

Then to find your possible rational zeros list all combinations of P/Q.

OpenStudy (anonymous):

So for my example the possible rational zeros are P/Q={1,2,4,8,1/5,2/5,4/5,8/5,-1,-2,-4, -8, -1/5,-2/5,-4/5,-8/5}. Now that is more than you would usually have but it is an example... If there are any repeats then you can toss the repeats out.

OpenStudy (anonymous):

i got -2,1,15 fir my answer

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

For your equation f(x) = x^3-4x^2-7x+10 ... What is your set P?

OpenStudy (anonymous):

Are you still with me @countrycuties16 ?

OpenStudy (anonymous):

yeah sorry i am completel lost

OpenStudy (anonymous):

completely*